OpenStudy (anonymous):

Which set of side lengths defines a right triangle? A. 6, 8, 11 B. 5, 12, 13 C. 7, 9, 13 D. 6, 9, 11

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

Do you know how to use the Pythagorean Theorem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, tell me then, what is the formula for the Pythagorean Theorem? That is what we need to know first to find out this answer.

OpenStudy (er.mohd.amir):

a^2+b^2=c^2 that is sum of square of two sides equal to square of third side

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

so then what

OpenStudy (anonymous):

Ok, remember that the hypotenuse is the longest side of a possible right triangle and is inserted in that formula as C while the other two sides are inserted in any order as A and B, try and plug in the sides keeping this in mind for the first possible answer, A.

OpenStudy (anonymous):

so 6^2+8^2=11^2 ?

OpenStudy (anonymous):

no its not A

OpenStudy (anonymous):

Right! Now try and repeat this for B.

OpenStudy (anonymous):

its B right?

OpenStudy (anonymous):

Try and use the pythagorean therom to find out.

OpenStudy (anonymous):

But yes, in doing so the answer would be B. Sorry I had to plug in my Laptop so it took me awhile to reply.